Total amount of 1,400 hundred dollars in two mutual funds. One earned 8% and the other earned 3% profit. The investor's profit was 67 dollars. How much was invested in each mutual fund.

Answer by checkley79(3341)

I'LL ASSUME THESE INVESTMENTS WERE FOR 1 YEAR.
.08X+.03(1400-X)=67
.08X+42-.03X=67
.05X=67-42
.05X=25
X=25/.05
X=$500 INVESTED @ 8%
1400-500=$900 INVESTED @ 3%
PROOF:
.08*500+.03*900=67
40+27=67
67=67